OpenStudy (anonymous):

x=-2-y,-5+y=-x-3 is this parallel, coincident, or intersecting

OpenStudy (anonymous):

seems to me like this would be parallel, but I don't know

OpenStudy (venomblast):

just set y='s if the slope is the same then it's parallel

OpenStudy (venomblast):

if not then set it equal to see if it's intersecting

OpenStudy (anonymous):

for -2-y I got 2y and for -5+y=-x-3 I got 2+xy , so I assumed they were parallel because they are basically the same.

OpenStudy (venomblast):

I could see this that if you set y=mx+b you would get y=-x-2 and y=-x+2

OpenStudy (venomblast):

these are two equations. How did you manage to multiply y and the x?

OpenStudy (anonymous):

I don't know honestly

OpenStudy (anonymous):

@StudyGurl14

OpenStudy (venomblast):